OpenStudy (anonymous):

Give the values of a, b, and c from the general form of the equation (2x + 1)(x - 2) = 0. Step by Step Please

Parth (parthkohli):

Y U NO MULTIPLY?

Parth (parthkohli):

\( \color{Black}{\Rightarrow (2x + 1) \times (x - 2) = 0 }\) After multiplying those,you'll get an equation in the form \(ax^2 + bx + c = 0\)

Parth (parthkohli):

Just compare the values of a, b and c with the equation you get.

Parth (parthkohli):

a would be the coefficient of \(x^2\). b would be the coefficient of \(x\) c would be the coefficient of \(x^0\) aka a constant.